OpenStudy (anonymous):

Property of exponent Simplify : x^8y^13/xy^6

jigglypuff314 (jigglypuff314):

property of exponents suggests that x^a/x^b = x^(a-b) so x^8/x^1 = x^(8-1) and y^13/y^6 = y^(13-6)

OpenStudy (anonymous):

is that the final answer

jigglypuff314 (jigglypuff314):

the code of conduct on this website advises us to not directly post answers for people :P your answer is the simplified of x^(8-1) y^(13-6) for the reasons exemplified above :)

OpenStudy (anonymous):

I understand the code,I just need help with it because I have already solved others problems .

jigglypuff314 (jigglypuff314):

x^8y^13/xy^6 = x^7 y^7